• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name V. Sulkosky

    Log entry time 15:04:28 on October 18, 2008

    Entry number 243996

    keyword=Left HRS VDC Spectra

    The left HRS has been re-connected to the BigBite DAQ as of Wednesday evening. Here are the first plots of the VDC spectra from run 1876 taken this morning.

    The VDC appears to be working fine.

    A copy of this log entry has been emailed to: bogdanw@jlab.org,segal@jlab.org,jiang@jlab.org



    Figure 1



    Figure 2



    Figure 3



    Figure 4